The mass of a block is 1kg.An electric heater of 48W takes 5 min to raise the temperature of the block from 20c to 50c.find the specific heat capacity of the block?
 

Grade:10

2 Answers

Pooja
4 years ago
Heat capacity = power × time ÷ temperature change 
For block of mass 1 kg
∆H = 48 × 5 × 60 ÷ 30 = 480 J/K
Khimraj
4 years ago
?H = power *time /t change inemperature = 48*5 *60 /30 = 480 J/K